OpenStudy (anonymous):

The Newmans family room is 16ft by 20ft. They are covering the floor with tiles that measure 1ft on a side and cost 0.82 each. How much would they spend on tile? @satellite73

OpenStudy (anonymous):

\[16\times 20\times .82\] is my guess

OpenStudy (anonymous):

390.24??

jimthompson5910 (jim_thompson5910):

The area is 16*20 = 320 square ft So you need 320/1 = 320 tiles Total Cost: 320*0.82 = $262.40